OpenStudy (anonymous):

HELP find critical points to a given problem dx/dt = x-y dy/dt = x^2 + y^2 -1 (phase plane analysis chapter DFQ)

OpenStudy (anonymous):

dy/dx= (x^2+y^2-1) / x- y dy/dx=0 implies (x^2+y^2-1) / x- y=0

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

wait then what do i do?

OpenStudy (anonymous):

do i integrate with respect to x?

OpenStudy (anonymous):

(x^2+y^2-1) / x- y=0 (x^2+y^2-1)=0

OpenStudy (anonymous):

so the critical point would be 0 ?

OpenStudy (anonymous):

have u got any values for x and y

OpenStudy (anonymous):

no. im really confused

OpenStudy (anonymous):

usually dx/dt and dy/dt equals to something in 't'. Have u got something like x=2t , y=1+t etc ???...otherwise you cant really go further than that

OpenStudy (anonymous):

everything i wrote at top is the only thing that it gave me

OpenStudy (anonymous):

ok then...here is the final word

OpenStudy (anonymous):

(x^2+y^2-1)=0 y^2=1- x^2 y= +- sqrt(1- x^2) are the critical points

OpenStudy (anonymous):

fan me if u liked my try to help u .thanx

OpenStudy (anonymous):

how do i fan you?